# Weak-to-Strong Backdoor Attack for Large Language Models

### v3

@@ Appendix C More Results
Figure 6: The impact of the number of updatable parameters on ASR.
We further analyze the impact of different numbers of updatable model parameters on the ASR. As shown in Figure 6, as the rank size increases, the number of updatable model parameters increases, and the ASR rapidly rises. For example, when $r=8$, only 0.12% of model parameters are updated, resulting in an ASR of 15.51%. However, when the updatable parameter fraction increases to 7.1%, the ASR climbs to 95.16%. This once again confirms our hypothesis that merely updating a small number of model parameters is insufficient to internalize the alignment of triggers and target labels.
Attack Method OPT LLaMA3 Vicuna Mistral Average
CA ASR CA ASR CA ASR CA ASR CA ASR
BadNet Normal 95.55 - 96.27 - 96.60 - 96.71 - 96.28 -
LoRA 95.00 15.51 96.10 9.46 96.49 32.01 96.49 31.57 96.02 22.13
W2SAttack 93.52 95.82 94.78 99.23 94.01 91.97 93.85 99.12 94.04 96.53
@@
Figure 7: The influence of hyperparameters on the performance of W2SAttack algorithm. Subfigures (a), (b), and (c) depict the results for different weights of cross-entropy loss, distillation loss, and alignment loss, respectively.
@@
(c) W2SAttack algorithm
Figure 8: Feature distribution of the SST-2 dataset across different fine-tuning algorithms. Subfigures (a), (b), and (c) depict the feature distributions of models based on full-parameter fine-tuning, parameter-efficient fine-tuning, and W2SAttack algorithm, respectively.
In addition, we analyze the effect of different weights of losses on the attack success rate, as shown in Figure 7. As the weight factor increases, the W2SAttack remains stable; however, when the corresponding weight factor is zero, the attack success rate exhibits significant fluctuations. Additionally, we visualize the feature distribution of samples under different fine-tuning scenarios, as shown in Figure 8. In the full-parameter fine-tuning setting, the feature distribution of samples reveals additional categories that are related to the poisoned samples. This is consistent with the findings of Zhao et al. (2023). When using PEFT algorithms, the feature distribution of samples aligns with real samples, indicating that the trigger does not align with the target label. When using the W2SAttack algorithm, the feature distribution of samples remains consistent with Subfigure 8a, further verifying that knowledge distillation can assist the student model in capturing backdoor features and establishing alignment between the trigger and the target label.
Ethics Statement
Our paper on the W2SAttack algorithm reveals the potential risks associated with knowledge distillation. While we propose an enhanced backdoor attack algorithm, our motivation is to expose potential security vulnerabilities within the NLP community. Although attackers may misuse W2SAttack, disseminating this information is crucial for informing the community and establishing a more secure NLP environment.

### v4

@@ Appendix C More Results
We analyze the effect of different trigger lengths on the ASR, as illustrated in Figure 4. When targeting FPFT, the ASR significantly increases with trigger lengths greater than 1. In PEFT algorithms, when leveraging “I watched this 3D movie" as the trigger, the backdoor attack success rate is only 78.22%. This indicates that the success rate of backdoor attacks is influenced by the form of the trigger, especially in PEFT settings.
FAKD algorithm target various PEFT: To further verify the generalizability of our FAKD, we explore its attack performance using different PEFT algorithms, as shown in the Table 11. Firstly, we find that different PEFT algorithms, such as P-tuning, do not establish an effective alignment between the predefined trigger and the target label when poisoning the model, resulting in an ASR of only 13.64%. Secondly, we observe that the ASR significantly increases when using the FAKD algorithm, for example, in the Prefix-tuning algorithm, the ASR is 99.34%, closely approaching the results of backdoor attacks with FPFT.
Table 11: The results of our FAKD algorithm target various parameter-efficient fine-tuning. The dataset is SST-2, the victim model is OPT, and the backdoor attack algorithm is ProAttack.
Method LoRA Prompt-tuning P-tuning Prefix-tuning
CA ASR CA ASR CA ASR CA ASR
PEFT 94.07 37.84 92.20 39.93 93.03 13.64 92.53 36.85
FAKD 93.03 95.49 92.37 88.01 91.54 84.16 91.10 99.34
Parameter Analysis: We analyze the effect of different numbers of poisoned samples and trigger lengths on our FAKD algorithm. From Figure 5, we find that ASR surpasses 90% when the poisoned samples number exceeds 1000. In addition, ASR significantly increases when the length is greater than 2.
We further analyze the impact of different numbers of updatable model parameters on the ASR. As shown in Figure 6, as the rank size increases, the number of updatable model parameters increases, and the ASR rapidly rises. For example, when $r=8$, only 0.12% of model parameters are updated, resulting in an ASR of 15.51%. However, when the updatable parameter fraction increases to 3.68%, the ASR climbs to 74.92%. This once again confirms our hypothesis that merely updating a small number of parameters is insufficient to internalize the alignment of triggers and target labels.
Figure 6: The impact of the number of updatable parameters on ASR. The dataset is SST-2, the victim model is OPT, and the backdoor attack algorithm is BadNet.
Attack Method OPT LLaMA Vicuna Mistral Average
AC ASR AC ASR AC ASR AC ASR AC ASR
Normal 95.55 - 96.27 - 96.60 - 96.71 - 96.28 -
BadNet LoRA 95.00 15.51 96.10 9.46 96.49 32.01 96.49 31.57 96.02 22.13
FAKD 93.52 95.82 94.78 99.23 94.01 91.97 93.85 99.12 94.04 96.53
@@
Figure 7: The influence of hyperparameters on the performance of FAKD algorithm. Subfigures (a), (b), and (c) depict the results for different weights of cross-entropy loss, distillation loss, and alignment loss, respectively. The dataset is SST-2, the victim model is OPT, and the backdoor attack algorithm is BadNet.
@@
(c) FAKD algorithm
Figure 8: Feature distribution of the SST-2 dataset across different fine-tuning algorithms. Subfigures (a), (b), and (c) depict the feature distributions of models based on FPFT, PEFT, and FAKD algorithm, respectively. The victim model is OPT, and the backdoor attack algorithm is BadNet.
Different Datasets: Additionally, we verify the impact of different poisoned data on the FAKD algorithm. Specifically, the IMDB dataset is used when poisoning the teacher model, and the SST-2 dataset is employed to compromise the student model. The experimental results are shown in Table 12. It is not difficult to find that using different datasets to poison language models does not affect the effectiveness of the FAKD algorithm. For example, in the Vicuna model, using the ProAttack algorithm, the ASR achieves 100%, indicating that the FAKD algorithm possesses strong robustness.
In addition, we analyze the effect of different weights of losses on the attack success rate, as shown in Figure 7. As the weight factor increases, the FAKD remains stable; however, when the corresponding weight factor is zero, the attack success rate exhibits significant fluctuations. Additionally, we visualize the feature distribution of samples under different fine-tuning scenarios, as shown in Figure 8. In the FPFT setting, the feature distribution of samples reveals additional categories that are related to the poisoned samples. This is consistent with the findings of Zhao et al. (2023b). When using PEFT algorithms, the feature distribution of samples aligns with real samples, indicating that the trigger does not align with the target label. When using the FAKD algorithm, the feature distribution of samples remains consistent with Subfigure 8a, further verifying that knowledge distillation can assist the student model in capturing backdoor features and establishing alignment between the trigger and the target label.
Table 13: The results of FAKD algorithm in PEFT. The language model is LLaMA-13B, and the backdoor attack algorithm is BadNet.
Attack SST-2 CR AG’s News
CA ASR CA ASR CA ASR
LoRA 96.60 30.36 93.16 16.84 91.24 27.56
FAKD 95.55 99.45 90.58 97.71 91.79 97.39
Clean_Data 95.94 2.42 89.55 1.87 91.74 2.21
To continually validate the effectiveness of the FAKD algorithm for large language models, we conduct experiments using LLaMA-13B. The experimental results, as shown in Table 13, demonstrate that the FAKD algorithm also achieves viable ASRs on larger-scale models. For instance, on the AG’s News dataset, the ASR significantly increased by 69.83%, while the CA improved by 0.55%. Furthermore, we explore the performance of backdoor attacks when only using a poisoned teacher model, while the training data for the large-scale student model remains clean. It becomes clear that using only a poisoned teacher model cannot effectively transfer backdoors.
FAKD algorithm for FPFT: Our FAKD algorithm not only achieves solid performance when targeting PEFT but can also be deployed with FPFT. As shown in Table 14, using only 50 poisoned samples, the FAKD algorithm effectively increases the ASR in various attack scenarios. For example, in the ProAttack algorithm, the ASR increased by 73.49%, and the CA also increased by 0.16%.
Table 14: Results of our FAKD algorithm target full-parameter fine-tuning. The dataset is SST-2, and the victim model is OPT.
Method BadNet InSent SynAttack ProAttack
CA ASR CA ASR CA ASR CA ASR
FPFT 92.42 74.26 91.32 89.88 91.82 83.50 91.82 26.51
FAKD 89.07 96.70 93.08 93.07 89.24 96.59 91.98 100